OpenStudy (goho2838):

G(s)=s^2+7 G(5)= G(-2)= G(0)=

OpenStudy (amorfide):

G(s)=s^2 + 7 this is a function of s if I had G(a) I would let s=a so G(a)=a^2 + 7 if I had g(4), I would let s=4 g(4)=4^2 + 7 so youwant G(5), so s=5 G(-2), so s=-2 G(0), so s=0

OpenStudy (amorfide):

can you solve this now|?

OpenStudy (goho2838):

Yes, thank you